In a shooting game the probabilities that kent and fred will hit a target are 2/3 and 3/4 respectively. What is the probability that the target is hit when both shoot at it?

In a shooting game the probabilities that kent and fred and mike will hit a target are 2/3 and 3/4 and 1/5 respectively. What is the probability that the target is hit when both shoot at it?

Answer by stanbon(48568)

In a shooting game the probabilities that kent and fred will hit a target are 2/3 and 3/4 respectively. What is the probability that the target is hit when both shoot at it?
P(both miss)=1/3 * 1/4 = 1/12
P(at least one will hit) = 1 - P(both miss)
= 1 - 1/12 = 11/12
============
In a shooting game the probabilities that kent and fred and mike will hit a target are 2/3 and 3/4 and 1/5 respectively. What is the probability that the target is hit when all of them shoot at it?
P(all miss)=1/3*1/4*4/5=1/15
P(at least one will hit) = 1 - P(all miss)
=1-1/15= 14/15
Cheers,
Stan H.